The length of a rectangle is 2 more than 5 times its width. Find the area of the rectangle if its perimeter is 64 in.
Answer by rothauserc(4718)

let l be the length of the rectange and w be the width of the rectangle
:
l = 5w +2
:
2l +2w = 64
:
2 * (5w+2) + 2w = 64
:
10w +4 +2w = 64
:
12w = 60
:
w = 5
:
l = 5 * 5 +2 = 27
:
Area of rectangle = 27 * 5 = 135 square inches
: